OpenStudy (anonymous):

Which of the following is not a rational number? 0.30030033 0/99 17 3/4 -100.1

OpenStudy (inkyvoyd):

just so you know, you may be missing symbols in this problem. Try using draw down below

OpenStudy (anonymous):

I think its B but.. I need help

OpenStudy (anonymous):

Imnot missing anything that's the whole question

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

No lines over any of the numbers?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

B is equal to 0, that's not it.

OpenStudy (anonymous):

Okay solo..

OpenStudy (anonymous):

I'm stuck now

OpenStudy (e.mccormick):

See, a rational number is anything that can be written completely as a ratio in numbers (no symbols). If I can do it as a ratio, it is rational. It includes all integers, so it includes representations of 0. \(0.30030033 = \dfrac{0.30030033}{1}= \dfrac{30030033}{100000000}\) so a ratio, and so on. http://www.mathsisfun.com/rational-numbers.html The 17 and 3/4 is a mixed number, but it can be made to an "improper" fraction and therefore a ratio.

OpenStudy (anonymous):

I'd go with A since it looks like 0.3003003300333003333.... whichis not rationsl. do any of the numbers have ANYTHING else ?

OpenStudy (anonymous):

No no numbers have anything else

OpenStudy (anonymous):

I say NONE

OpenStudy (anonymous):

So it would be c? @e.mccormick

OpenStudy (e.mccormick):

Everthing there, as shown, could be written as a rational number.

OpenStudy (e.mccormick):

If it was 99/0, that would not be rational. If \(e\), \(i\), or \(\pi\) was there, it would not be rational.

OpenStudy (anonymous):

A has three dots written after it

OpenStudy (e.mccormick):

AH HA! See, that means it continues!

OpenStudy (e.mccormick):

If it continues it is not rational.

OpenStudy (anonymous):

So A!! OMG THANKS!

OpenStudy (e.mccormick):

Yah, that is why inkyvoyd, Allofus, and I all said it looked like something was missing. Hehe.

OpenStudy (anonymous):

Yeah I'm sorry their was SOMETHING heir sorry guys and girls

OpenStudy (e.mccormick):

Well, it got found in the end and now you know what it means, so all good! That is the learning process.

OpenStudy (anonymous):

Could you help with more?

OpenStudy (e.mccormick):

Just close this and make a new question. Plenty of people to help. I am going to have to get back to something else soon.